You lent $370 to a friend for one year at a nominal rate of interest of 3 percent. Inflation during that year was 2 percent. Did you experience an increase or decrease in the purchasing power of your money? How much did it increase or decrease? (Round answer to 2 decimal places, e.g. 52.75%.)

The purchasing power ______(increasing or decreasing) by ____%.

If the nominal rate of interest is 4.19 percent and the expected rate of inflation is 1.76 percent, what is the real rate of interest? (Round percentage answer to 3 decimal places, e.g. 52.754%.)

real rate of interest ___%
Answer #1

1.

Real Growth Rate = 0.03 - 0.02 = 1.00%

The purchasing power increases by 1%

2.

Real Rate = (1.0419/1.0176) - 1

Real Rate = 2.388%